A mathematics class consist of 15 boys and 25 girls. A.In how many ways can the position of President, Vise-President, secretary, treasurer, and auditor be filled if no student may hold more than one position? B.In how many ways can 5 representatives be chosen from the class? C.If exactly 2 girls are chosen, in how many ways can the 5 representatives be chosen?
Answer by stanbon(75887)

A mathematics class consist of 15 boys and 25 girls.
A.In how many ways can the position of President, Vise-President, secretary, treasurer, and auditor be filled if no student may hold more than one position?
-------
Ans: 50P5 = 50*49*48*47*46
--------------------------------------
B.In how many ways can 5 representatives be chosen from the class?
Ans: 50C5 = 2118760
--------------------------------------
C.If exactly 2 girls are chosen, in how many ways can the 5 representatives be chosen?
Ans: 25C2*15C3 = 300*455 = 136,500
=======================================
Cheers,
Stan H.